# CoA: Chain-of-Action for Generative Semantic Labels

### v1

@@ Abstract
Recent advances in vision-language models (VLM) have demonstrated remarkable capability in image classification. These VLMs leverage a predefined set of categories to construct text prompts for zero-shot reasoning. However, in more open-ended domains like autonomous driving, using a predefined set of labels becomes impractical, as the semantic label space is unknown and constantly evolving. Additionally, fixed embedding text prompts often tend to predict a single label (while in reality, multiple labels commonly exist per image). In this paper, we introduce CoA, an innovative Chain-of-Action (CoA) method that generates labels aligned with all contextually relevant features of an image. CoA is designed based on the observation that enriched and valuable contextual information improves generative performance during inference. Traditional vision-language models tend to output singular and redundant responses. Therefore, we employ a tailored CoA to alleviate this problem. We first break down the generative labeling task into detailed actions and construct an CoA leading to the final generative objective. Each action extracts and merges key information from the previous action and passes the enriched information as context to the next action, ultimately improving the VLM in generating comprehensive and accurate semantic labels. We assess the effectiveness of CoA through comprehensive evaluations on widely-used benchmark datasets and the results demonstrate significant improvements across key performance metrics. Our source code is available at: https://github.com/WilsonMqz/CoA
1 Introduction
Visual-language models have achieved significant success in various downstream tasks encompassing both vision and language, particularly in image classification tasks [3, 18, 16, 36]. These models demonstrate remarkable results in zero-shot image classification tasks by aligning the semantics of vision and language [34, 23, 24, 29]. However, this alignment relies on a predefined set of categories (as illustrated on the left of Fig. 1), achieved by constructing embedding templates such as “a photo of a class” for zero-shot reasoning [34, 23]. This assumption becomes impractical with the increasingly vast real-world data scale where semantic label spaces are often unknown and continuously evolving [8, 50, 27]. For instance, in domains like autonomous driving and robot vision, data collected in complex environments is diverse, while label spaces dynamically evolve over time [42, 13, 14, 25].
In this work, we consider removing the assumption of a predefined set of labels and focus on directly assigning semantic-level labels to an image in a more unbounded environment, known as Generative Semantic Labels (GSLs). It is worth noting that the problem we address differs significantly from the traditional open vocabulary problem, which assumes the model has knowledge of and access to the list of target classes during inference [34, 23, 24]. Coinciding with our approach, Conti et al. [8] introduced Vocabulary-free Image Classification (VIC), which aims to assign a single label to an image without the assumption of predefined set of categories. However, they need to search through a vast visual-language library to find a similar semantic sentence for the image, essentially providing a larger semantic space. On the other hand, VIC’s objective is to match every image with one label, which is unrealistic in practical downstream tasks like semantic segmentation and object detection, where each image often has multiple labels (as depicted on the center of Fig. 1). Therefore, this inspired us to investigate a novel approach capable of predicting all labels within an image without assumption for a predefined semantic space.
Figure 1: Comparison of the proposed method with other methods. (a) Vision-Language Model (VLM)-based methods operate under the assumption of a predefined set of target categories and typically output a single label per image. (b) CaSED removes the constraint of predefined categories but requires retrieval from an extensive text database and still ultimately predicts only a single label. (c) In contrast, our proposed Chain-of-Action (CoA) method simultaneously predicts multiple labels for an image without the need for predefined categories or retrieval, offering a more comprehensive and direct approach to semantic-level classification.
In recent years, Large Multimodal Models (LMMs) such as LLaVA [29], GPT-4V [1], and InstructBLIP [9] have showcased impressive advancements in the realm of vision and language (VL), particularly in multimodal reasoning and visual question-answering (VQA). Therefore, we naturally consider two straightforward approaches to handle the GSLs task: VQA-based LMMs, which directly ask questions to obtain entity labels, and Caption-based LMMs, which first generate captions and then filter relevant entity labels. However, as shown in Fig. 2, VQA-based methods produce semantically overlapping labels, while Caption-based methods tend to generate background or extended semantic labels—both of which are not well-suited for GSLs. Further discussion and analysis are provided in Section 3.2.
In this paper, we introduce a zero-shot, free vocabulary Chain-of-Action (CoA) method to better handle GSls task. Our approach is designed on the observation that enriching the inference process with valuable contextual information leads to enhanced model performance [31, 46]. This capability of context-based prompting has also been widely validated in the application of large language models [21, 44].
Figure 2: Comparison of various settings, including VQA-based, Caption-based, Human-labeled and the proposed CoA method. VQA-based approaches often suffer from significant semantic redundancy, while Caption-based methods tend to identify background-related or extended semantics. In our evaluation, the dataset-provided labels are used as human annotations, which are relatively limited. Areas that are ignored are highlighted and the proposed CoA method generates results that are more comprehensive, diverse, and accurate.
In our work, we define a single interaction with LMMs as an individual action and design a progressive hierarchical chain-of-action structure to guide LMMs in generating the semantic labels. Initially, we leverage LMMs to produce an overall description of the test image, extracting and consolidating key local information. We then refines the representation of the key entities and their interrelationships, using this structured data as context. This progressive structure enriches the contextual understanding of LMMs, ultimately facilitating more comprehensive and precise label prediction. Our approach aligns with the principles of In-Context Learning by enhancing generative performance through contextual utilization but diverges by eliminating the need for similar, unrelated examples or additional data. Instead, it utilizes the inherent capabilities of LMMs and is adaptable across various models.
To quantitatively assess the performance of methods addressing GSLs task, we introduced a set of metrics to measure how effectively the predicted semantic labels align with the test image. Across all tasks and metrics, our methods consistently outperform all baselines, including sophisticated VLMs such as BLIP-2 and Instruct-BLIP. It is worth mentioning that our approach is simple yet effective and can serve as a competitive baseline for future research endeavors aimed at addressing the challenging GSLs task.
In summary, our main contributions are as follows:
@@ 1 Introduction
(1) We explore a new task of generating semantic-level labels in a more open environment, eliminating the assumption of predefined label sets in VLMs. We formalize this task and propose specific evaluation metrics that can serve as a guideline for future studies.
@@
(2) We introduce a customized zero-shot Chain-of-Action (CoA) approach that effectively guides LMMs in producing comprehensive and accurate semantic labels. Additionally, our CoA method is adaptable across various LMMs architectures.
@@
2 Related Work
2.1 Large MultiModal Models (LMMs)
Recent advancements in Large Multimodal Models (LMMs) [2, 43, 9, 10, 24, 28, 29, 57] have successfully integrated the sophisticated reasoning abilities [35, 38] of large language models with visual-language models (VLM) [16, 23, 34], leading to significant breakthroughs in a wide range of downstream tasks. These tasks include, but are not limited to, Visual Question Answering (VQA) [3, 18, 30, 17, 19, 36], image-to-text generation, and text-to-image synthesis. Current approaches in LLMs predominantly focus on two core strategies: the alignment of image-text features and instruction fine-tuning.
Image-Text Feature Alignment. The image-text feature alignment approach is exemplified by CLIP [34], a pioneering work in the multimodal domain. CLIP employs 400 million image-text pairs to align text and image features within a shared embedding space. Building on this foundation, BLIP [23] integrates visual-text comprehension and generation within a unified multimodal framework, utilizing joint pre-training on three key objectives: image-text contrastive learning, image-text matching, and image-conditioned language modeling. Further advancing this line of work, BLIP-2 [24] capitalizes on pre-trained VLM to enhance multimodal performance and streamline the training process.
Instruction Fine-Tuning. In the instruction fine-tuning paradigm, InstructBLIP [9] integrates instruction-based inputs to guide both the Q-former and the large language model (LLM). Mini-GPT4 [58] employs a projection layer to bridge a visual encoder with a more advanced LLM, focusing training exclusively on this projection layer. Similarly, LLaVA [29] leverages GPT-4 [1] to generate multimodal instruction-following vision-test data and connects the visual encoder to the LLM via a projection layer, which is trained using an autoregressive model loss.
These approaches predominantly explore contrastive learning techniques to investigate refine graph-text feature alignment and enhance instruction fine-tuning capabilities. Although they have achieved significant success in complex tasks related to image-text comprehension and generation, these methods exhibit limitations in handling simpler classification tasks. One-step generative answers are often redundant, primarily due to the rich semantics embedded in the image-text training data of LLMs. In contrast, our approach utilizes a step-by-step process to guide LLMs in generating more comprehensive and accurate information.
2.2 Large language models / Chain-of-Thought
@@
Large language models (LLMs) [33, 35, 6, 38, 51, 39, 11, 49] have garnered significant attention in recent years due to their ability to understand and generate natural language text. Models like GPT-3 [6] and LLaMA [39] have demonstrated impressive performance across various natural language processing tasks. These models are typically pre-trained on massive text data and fine-tuned on specific downstream tasks [39, 11, 49]. They have been employed in applications such as text generation, translation, summarization, and question answering. Research efforts have focused on enhancing the capabilities, efficiency, and interpretability of these large language models.
@@
The chain-of-thought (CoT) [45, 56, 53, 41, 47, 22, 48, 55, 5, 40, 54] technique has emerged as a valuable approach to guide and structure the generation of content by LLMs. By utilizing a progressive hierarchical chain structure, researchers aim to enhance the coherence and contextuality of model-generated text [45, 56, 53]. This technique serves as a mechanism to direct the flow of information and ensure the generation of relevant and meaningful content [47, 22]. CoT play a crucial role in constraining the output of LLMs, enabling them to produce text that aligns closely with desired contexts and themes. Integrating the CoT technique with LLMs has the potential to improve the quality and relevance of generated text across a wide range of applications and use cases.
@@
3 Method
Figure 3: Pipeline of CoA. First, we utilize Caption Action to generate a comprehensive image description and extract an initial object list. Self-Correct Action refines this list, followed by Appearance Action and Relationship Action to capture detailed appearance features and relationships. Finally, we integrate these responses to create enriched context, guiding LMMs to achieve final semantic labels.
@@ 3.1 Preliminaries
VLM-based Image Classification. Let $\mathcal{X}\subset\mathbb{R}^{d}$ represent the $d$-dimensional image space, and $\mathcal{Y}=\{1,2,\ldots,k\}$ denote the label space with $k$ distinct classes. VLM-based methods construct text prompts in the form of templates, such as “a photo of a $\{CLASS\}$”, where $CLASS$ is a placeholder for an element from $\mathcal{Y}$ [34]. The text encoder in the VLM processes these templates to generate text embeddings, denoted as $T$. Simultaneously, for a given input image, the visual encoder maps the image into its corresponding visual embedding vector $V$. The core of VLM-based image classification lies in aligning the semantic spaces of $V$ and $T$. VLM-based methods are typically trained on large-scale image-text datasets to learn a mapping function $f:\mathcal{X}\mapsto\mathcal{Y}$, where the objective is to match image representations to their corresponding textual labels. However, in standard VLM-based image classification, the label space $\mathcal{Y}$ is manually pre-defined and fixed, which becomes impractical in the real-world scenarios where $\mathcal{Y}$ is unknown and subject to continuous change.
@@
Vocabulary-free Image Classification. Given the image space $\mathcal{X}$, Vocabulary-free Image Classification (VIC) aims to assign a class label $y$ to an test image $x$ without prior knowledge of the label set $\mathcal{Y}$ [8]. Instead, VIC assumes access to a broader semantic class space $\mathcal{S}$ (where $\mathcal{Y}\subset\mathcal{S}$), which encompasses all potential semantic concepts. The objective of VIC is to define a function $f$ that maps an image to a semantic label within $\mathcal{S}$, i.e., $f:\mathcal{X}\mapsto\mathcal{S}$. However, the vastness of this semantic space presents a significant challenge, particularly when distinguishing fine-grained concepts across diverse domains or handling long-tailed distributions. As such, this assumption often proves impractical in real-world applications.
3.2 Generative semantic labels
@@
Task Definition. Given a test image $x_{i}\in\mathcal{X}$, Generative Semantic Labels (GSLs) aims to assign a semantic labels set $Y_{i}$ to the test image $x_{i}$ without prior knowledge of the label set $\mathcal{Y}$. Here, $Y_{i}\in\mathcal{C}$ is the set of relevant semantic labels associated with $x_{i}$, where $\mathcal{C}=2^{\mathcal{S}}$. The goal of GSLs is to induce a map function $f:\mathcal{X}\mapsto 2^{\mathcal{S}}$. Notably, the key distinction between GSLs and VIC lies in the fact that GSLs assign multiple semantic labels to the test image $x_{i}$, rather than a single class label. Moreover, the VIC task requires additional methods to obtain a large semantic space $\mathcal{S}$ and to retrieve the most relevant semantic label for a given test image. In contrast, GSLs does not rely on such assumptions, as $S$ is inherently the knowledge base of the LMMs, with the semantics derived solely from the model’s internal knowledge.
@@
The primary challenge of the GSLs task lies in extracting the semantics of all entities within the test image. The core issue in method design is how to accomplish this without relying on predefined labes set or external semantic databases. Our only resource is the generalized knowledge embedded in LMMs. Therefore, fully leveraging the prior knowledge learned by these models during training becomes the most critical problem to solve in GSLs task. To address this, we can either directly extract entity semantics using the model’s image understanding capabilities or generate detailed textual descriptions via the captioning abilities of LMMs, which can then be parsed for entity semantics. Based on this, we propose two straightforward baseline approaches: a single-stage method, using VQA-based LMMs to directly query the model for entity semantics, and a two-stage method, where captions are generated and then filtered to identify relevant entities, referred to as Caption-based LMMs.
@@
VQA-based LMMs. Large Multimodal Models (LMMs) are powerful tools with human-competitive image understanding capabilities, capable of handling complex visual question-answering tasks. In this paper, we leverage this ability to extract high-quality entity semantics by fine-tuning LLaVA [29], a popular multimodal large model. We design various question templates to guide the model in generating the desired outputs (The detailed information regarding these question templates can be found in Appendix A). However, we observed that directly querying the model often yields suboptimal results. Specifically, the confidence scores for predicted labels tend to favor entities with particularly prominent features. Moreover, the generated labels frequently exhibit high semantic overlap, which is not aligned with the desired outcomes, as illustrated in Fig. 2.
@@
Caption-based LMMs. To address the issue of semantic redundancy in VQA-based methods, we explore a captioning approach for the GSLs task. This involves first generating detailed image descriptions using LMMs, followed by filtering the captions to obtain the final set of entity semantics. The generated captions often capture fine-grained details, resulting in more comprehensive and diverse outputs. However, we observed that excessively long captions incur substantial computational cost, while shorter captions may overlook certain labels and ignore local details. This process requires significant parameter tuning to achieve optimal results. In addition, this approach is also constrained by the limitations of the filtering techniques. Since the image descriptions are generated based on the input image, the LMMs’ associative capabilities can lead to hallucinations, where the output includes entity semantics that do not accurately match the image content. As illustrated in Fig. 2, inadequate filtering may exacerbate this issue.
3.3 Chain-of-Action
@@
Figure 4: The instruction templates defined within the various action setting.

### v2

# Seeing the Undefined: Chain-of-Action for Generative Semantic Labels

@@ Abstract.
Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in image classification by leveraging predefined sets of labels to construct text prompts for zero-shot reasoning. However, these approaches face significant limitations in undefined domains, where the label space is vocabulary-unknown and composite. We thus introduce Generative Semantic Labels (GSLs), a novel task that aims to predict a comprehensive set of semantic labels for an image without being constrained by a predefined labels set. Unlike traditional zero-shot classification, GSLs generates multiple semantic-level labels, encompassing objects, scenes, attributes, and relationships, thereby providing a richer and more accurate representation of image content. In this paper, we propose Chain-of-Action (CoA), an innovative method designed to tackle the GSLs task. CoA is motivated by the observation that enriched contextual information significantly improves generative performance during inference. Specifically, CoA decomposes the GSLs task into a sequence of detailed actions. Each action extracts and merges key information from the previous step, passing enriched context to the next, ultimately guiding the VLM to generate comprehensive and accurate semantic labels. We evaluate the effectiveness of CoA through extensive experiments on widely-used benchmark datasets. The results demonstrate significant improvements across key performance metrics, validating the capability of CoA to generate accurate and contextually rich semantic labels. Our work not only advances the state-of-the-art in generative semantic labels but also opens new avenues for applying VLMs in open-ended and dynamic real-world scenarios. Our source code is available at: https://github.com/WilsonMqz/CoA
Generative semantic labels, Chain-of-Action, Vision-language model, Vocabulary-Unknown, Composite labels
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755157isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Object recognition
1. Introduction
[…]
However, a significant challenge lies in the fact that predefining a set of labels is often impractical in real-world undefined environments. The label space in real-world scenarios is typically vocabulary-unknown and composite, which makes traditional pre-defined-label-set approaches inadequate (Conti et al., 2023; Zhang et al., 2024a; Liu et al., 2023). Especially in complex scenes, the semantic information of many objects or concepts is composite and cannot be simply represented by a single label. For example, a composite label like ”humanoid landmark” encompasses both visual features (humanoid) and functional attributes (landmark), a composite semantic that neither exists in traditional predefined labels set nor can be accurately described by a single category label. Similarly, in the field of medical imaging, newly discovered pathologies (such as variants of COVID-19) or rare conditions require models to adapt to an ever-expanding labels space, where these new categories are often vocabulary-unknown, as illustrated in Fig. 1.
To address this challenge, we formalize Generative Semantic Labels (GSLs), a novel task that eliminates the need for predefined labels set and empowers models to directly assign a comprehensive set of semantic-level labels to an image. Unlike traditional open-vocabulary classification, which relies on access to a predefined labels set during inference (Radford et al., 2021; Li et al., 2022, 2023), GSLs operates in a fully undefined environment, generating labels that capture the richness and diversity of real-world semantics. Furthermore, GSLs predicts multiple and composite labels per image, aligning with practical downstream tasks such as semantic segmentation and object recognition, where images inherently contain multiple semantic elements (more details can be found in Sec. 3). The advantages of GSLs are twofold: (1) flexibility—unconstrained by predefined labels set, GSLs adapts to diverse and undefined domains; and (2) comprehensiveness—it generates multiple composite semantic labels, capturing the full context of an image.
To tackle the GSLs task, we propose Chain-of-Action (CoA), a zero-shot, generative method that leverages the contextual reasoning capabilities of Vision-Language Models (VLMs) like LLaVA (Liu et al., 2024c) and InstructBLIP (Dai et al., 2023). In our study, CoA decomposes the generative process into a sequence of progressive actions. Each action extracts and refines key contextual information, enabling the model to generate accurate and comprehensive semantic labels. Specifically, CoA first generates an overall description of an image, then identifies key entities and their relationships, and finally consolidates these information into a structured set of labels. This hierarchical approach aligns with the principles of in-context learning (Mitra et al., 2024; Yao et al., 2024a) but eliminates the need for external examples or additional data, making it both efficient and adaptable.
Figure 1. An example of the Generative Semantic Labels (GSLs) task. In real-world scenarios, the label space is often vocabulary-unknown and composite. Taking the COVID-19 and its variants as an example, a predefined labels set does not include labels for new variants such as ”Omicron” which limits the recognition capabilities of visual-language models (VLM). The GSL task removes the assumption of a predefined labels set, enabling the identification of undefined labels and generating composite semantic labels related to the image.
We evaluate CoA on widely-used benchmarks, introducing a set of metrics to measure the alignment between predicted labels and ground truth. Our results demonstrate consistent improvements over state-of-the-art baselines, including BLIP-2 and InstructBLIP, across all tasks and metrics. Notably, CoA’s simplicity and effectiveness make it a strong baseline for future research on GSLs. By enabling models to generate comprehensive, dynamic, and contextually rich semantic labels, our work paves the way for more robust and adaptable vision-language systems in open-ended domains. In summary, our main contributions are as follows:
@@ 1. Introduction
(1) We explore a new task of generating semantic-level labels in a more open environment, eliminating the assumption of predefined labels set in VLMs. We formalize this task and propose specific evaluation metrics that can serve as a guideline for future studies.
@@
(2) We introduce a customized zero-shot Chain-of-Action (CoA) approach that effectively guides VLMs in producing comprehensive and accurate semantic labels. Additionally, our CoA method is adaptable across various VLMs architectures.
@@
2. Related Work
2.1. Open-Vocabulary Image Classification
Recent advancements in Vision-Language Models (VLMs) (Alayrac et al., 2022; Wei et al., 2022a; Dai et al., 2023; Driess et al., 2023; Li et al., 2023; Liu et al., 2024a, c; Zhu et al., 2024a; Goel et al., 2022; Li et al., 2022; Radford et al., 2021) have combined the reasoning capabilities of large language models (Raffel et al., 2020; Tay et al., 2023) with visual information, achieving significant breakthroughs in zero-shot open-vocabulary image classification through two core strategies: image-text feature alignment (Radford et al., 2021; Li et al., 2022, 2023) and instruction fine-tuning (Dai et al., 2023; Liu et al., 2024c; Zhu et al., 2024b). Image-text alignment, exemplified by CLIP (Radford et al., 2021) and extended by BLIP (Li et al., 2022) and BLIP-2 (Li et al., 2023), aligns visual and textual features in a shared embedding space. Instruction fine-tuning, as seen in InstructBLIP (Dai et al., 2023), Mini-GPT4 (Zhu et al., 2024b), and LLaVA (Liu et al., 2024c), leverages instruction-based inputs and projection layers to connect visual encoders with advanced LLMs, optimizing multimodal performance.
[…]
2.2. Vocabulary-Free Image Classification
Recently, a new paradigm has emerged, known as vocabulary-free image classification (VIC) (Conti et al., 2023; Zhang et al., 2024a; Liu et al., 2023), which aims to assign a single label to an image without the assumption of predefined set of categories. However, they need to search through a vast visual-language library to find a similar semantic sentence for the image (Conti et al., 2023), essentially providing a larger semantic labels space. On the other hand, VIC’s objective is to match each image with one label, which is unrealistic in practical downstream tasks like semantic segmentation and object detection, where each image often has multiple labels. Moreover, the labels generated by VIC carry single semantic information, which cannot fully capture the content of an image. In contrast, GSLs can generate multiple composite semantic labels for each image.
2.3. Large language models / Chain-of-Thought
@@
Large language models (LLMs) (Radford et al., 2019; Raffel et al., 2020; Brown et al., 2020; Tay et al., 2023; Zhang et al., 2022; Touvron et al., 2023; Du et al., 2022; Zeng et al., 2023) have garnered significant attention in recent years due to their ability to understand and generate natural language text. Models like GPT-3 (Brown et al., 2020) and LLaMA (Touvron et al., 2023) have demonstrated impressive performance across various natural language processing tasks. These models are typically pre-trained on massive text data and fine-tuned on specific downstream tasks (Touvron et al., 2023; Du et al., 2022; Zeng et al., 2023). They have been employed in applications such as text generation, translation, summarization, and question answering. Research efforts have focused on enhancing the capabilities, efficiency, and interpretability of these large language models.
Figure 2. Pipeline of CoA. First, we utilize Caption Action to generate a comprehensive image description and extract an initial object list. Self-Correct Action refines this list, followed by Appearance Action and Relationship Action to capture detailed appearance features and relationships.  Finally, we integrate these responses to create enriched context, guiding VLMs to achieve final semantic labels.
@@
The chain-of-thought (CoT) (Wei et al., 2022b; Zhou et al., 2023; Zhang et al., 2023a; Wang et al., 2023; Yao et al., 2024b; Lei et al., 2023; Yao et al., 2023; Zheng et al., 2023; Besta et al., 2024; Wang et al., 2024; Zhang et al., 2024c) technique has emerged as a valuable approach to guide and structure the generation of content by LLMs. By utilizing a progressive hierarchical chain structure, researchers aim to enhance the coherence and contextuality of model-generated text (Wei et al., 2022b; Zhou et al., 2023; Zhang et al., 2023a). This technique serves as a mechanism to direct the flow of information and ensure the generation of relevant and meaningful content (Yao et al., 2024b; Lei et al., 2023). CoT play a crucial role in constraining the output of LLMs, enabling them to produce text that aligns closely with desired contexts and themes. Integrating the CoT technique with LLMs has the potential to improve the quality and relevance of generated text across a wide range of applications and use cases.
@@
3. Method
@@ 3.1. Preliminaries
Open-Vocabulary Image Classification. Let $\mathcal{X}\subset\mathbb{R}^{d}$ represent the $d$-dimensional image space, and $\mathcal{Y}=\{1,2,\ldots,k\}$ denote the label space with $k$ distinct classes. Open-Vocabulary methods construct text prompts in the form of templates, such as ”a photo of a $\{CLASS\}$”, where $CLASS$ is a placeholder for an element from $\mathcal{Y}$ (Radford et al., 2021). Open-Vocabulary methods are typically trained on large-scale image-text datasets to learn a mapping function $f:\mathcal{X}\mapsto\mathcal{Y}$, where the objective is to match image representations to their corresponding textual labels. However, in standard Open-Vocabulary image classification, the label space $\mathcal{Y}$ is manually predefined and fixed, which becomes impractical in the real-world scenarios where $\mathcal{Y}$ is unknown and subject to continuous change.
@@
Vocabulary-free Image Classification. Similarly, VIC assumes access to a broader semantic class space $\mathcal{S}$ (where $\mathcal{Y}\subset\mathcal{S}$), which encompasses all potential semantic concepts. The objective of VIC is to define a function $f$ that maps an image to a semantic label within $\mathcal{S}$, i.e., $f:\mathcal{X}\mapsto\mathcal{S}$. However, the vastness of this semantic space presents a significant challenge, particularly when distinguishing fine-grained concepts across diverse domains or handling long-tailed distributions. As such, this assumption often proves impractical in real-world applications.
3.2. Generative Semantic Labels
@@
Task Definition. Given a test image $x_{i}\in\mathcal{X}$, Generative Semantic Labels (GSLs) aims to assign a semantic labels set $Y_{i}$ to the test image $x_{i}$ without prior knowledge of the label set $\mathcal{Y}$. Here, $Y_{i}\in\mathcal{C}$ is the set of relevant semantic labels associated with $x_{i}$, where $\mathcal{C}=2^{\mathcal{S}}$. Notably, the key distinction between GSLs and VIC lies in the fact that GSLs assign multiple semantic labels to the test image $x_{i}$, rather than a single class label. Moreover, the VIC task requires additional methods to obtain a large semantic space $\mathcal{S}$ and to retrieve the most relevant semantic label for a given test image. In contrast, GSLs does not rely on such assumptions, as $S$ is inherently the knowledge base of the VLMs, with the semantics derived solely from the model’s internal knowledge.
Seeing the Undefined. As shown in Fig. 1, the proposed GSLs is capable of updating in real-time in response to changes over time and user needs, not being confined to outdated, manually predefined labels set. Secondly, GSLs has the ability to ”see the undefined”, that is, it can recognize unknown vocabulary not covered by the fixed manually-predefined labels set, making it more adaptable to the complex and diverse tasks and scenarios of the real world, such as in the fields of novel disease detection and content moderation.
GSLs vs multi-label system. Conventional multi-label learning tasks are limited in two key aspects: they are unable to discover unknown categories and fail to represent complex composite concepts. For example, the COCO dataset often contains images depicting ”Five people and two boats.” While existing multi-label systems can detect the presence of ”people” and ”boat” as separate labels, they are incapable of capturing the compositional semantics of ”five people near two boats,” which better align with real-world requirements. As illustrated in Fig. 1, GSLs introduces composite semantic labels through CoA, addressing both open-vocabulary detection and semantic compositionality.
@@
How to Tackle GSLs. The primary challenge of the GSLs task lies in extracting the semantics of all entities within the test image. The core issue in method design is how to accomplish this without relying on predefined labels set or external semantic databases. Our only resource is the generalized knowledge embedded in VLMs. Therefore, fully leveraging the prior knowledge learned by these models during training becomes the most critical problem to solve in GSLs task. To address this, we can either directly extract entity semantics using the model’s image understanding capabilities or generate detailed textual descriptions via the captioning abilities of VLMs, which can then be parsed for entity semantics. Based on this, we propose two straightforward baseline approaches: a single-stage method, using VQA-based VLMs to directly query the model for entity semantics, and a two-stage method, where captions are generated and then filtered to identify relevant entities, referred to as Caption-based methods.
@@
VQA-based methods. VLMs are powerful tools with human-comp-etitive image understanding capabilities, capable of handling complex visual question-answering tasks. In this paper, we leverage this ability to extract high-quality entity semantics by fine-tuning LLaVA (Liu et al., 2024c), a popular multimodal large model. We design various question templates to guide the model in generating the desired outputs. However, we observed that directly querying the model often yields suboptimal results. Specifically, the confidence scores for predicted labels tend to favor entities with particularly prominent features. Moreover, the generated labels frequently exhibit high semantic overlap, which is not aligned with the desired outcomes, as illustrated in Fig. 3.
@@
Caption-based methods. To address the issue of semantic redundancy in VQA-based methods, we explore a captioning approach for the GSLs task. This involves first generating detailed image descriptions using VLMs, followed by filtering the captions to obtain the final set of entity semantics. The generated captions often capture fine-grained details, resulting in more comprehensive and diverse outputs. However, we observed that excessively long captions incur substantial computational cost, while shorter captions may overlook certain labels and ignore local details. This process requires significant parameter tuning to achieve optimal results. In addition, this approach is also constrained by the limitations of the filtering techniques. Since the image descriptions are generated based on the input image, the VLMs’ associative capabilities can lead to hallucinations, where the output includes entity semantics that do not accurately match the image content. As illustrated in Fig. 3, inadequate filtering may exacerbate this issue.
Figure 3. Comparison of various settings, including VQA-based, Caption-based, Human-labeled and the proposed CoA method. In our evaluation, the dataset-provided labels are used as human annotations, which are relatively limited. Areas that are ignored are highlighted and the proposed CoA method generates results that are more comprehensive, diverse, and accurate.
@@
Figure 4. The instruction templates defined within the various action setting.
3.3. Chain-of-Action